OpenStudy (anonymous):

Solve. 14y + 20 = 4y A. –2 B. 2 C.no solution D.all real numbers

OpenStudy (anonymous):

A. -2

OpenStudy (anonymous):

BECAUSE

OpenStudy (anonymous):

So first want to get the like terms on the same side of the equation. -20=10y. Divide by 10 and y=-2

OpenStudy (anonymous):

tep 1 : Pulling out like terms : 1.1 Pull out like factors: 10y + 20 = 10•(y + 2) Equation at the end of step 1 : 10 • (y + 2) = 0 Step 2 : Solve 10•(y+2) = 0 Equations which are never true : 2.1 Solve 10 = 0 This equation has no solution(s). A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : y+2 = 0 Subtract 2 from both sides of the equation: y = -2

OpenStudy (anonymous):

Done.

OpenStudy (anonymous):

**Step

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np